OpenStudy (anonymous):

evaluate the function at the given values of x, if possible f(x)= 2x-1/x^2 (a) x=2 (b) x=-1

OpenStudy (anonymous):

a) 3/4 b)-3 did u really not kno how to do that?

OpenStudy (anonymous):

a f(x)= 2(2)-1/(2)^2=15/4 b f(x)= 2(-1)-1/(-1)^2=-3

OpenStudy (anonymous):

\[f(x)=\frac{2x-1}{x^2}\] \[f(2)=\frac{2\times 2-1}{2^2}=\frac{4-1}{4}=\frac{3}{4}\]

OpenStudy (anonymous):

\[f(-1)=\frac{2\times -1-1}{(-1)^2}=\frac{-2-1}{1}=-3\]

OpenStudy (anonymous):

the way its written a) is 15/4 but i think he meant to write it the way satellite73 wrote it